OpenStudy (anonymous):

Which of the following points lie in the solution set to the following system of inequalities? y < -3x + 3 y < x + 2 (1, -5) (1, 5) (5, 1) (-1, 5)

OpenStudy (anonymous):

substitute each point and check a)5,2 in y<=x-5 2<=5-5 => 2<=0 this is not true so this is not the answer b) 5,-2 in -2<=5-5 -2<=5-5 -2<=0 this is true let's check with the other inequality y>=-x-4 -2>=-5-4 true so option B is one of the answers c) (-5,-2) y<=x-5 -2<=-5-5 -2<=-10 false d)(-5,2) 2<=-5-5 2<=-10 this is also not true SO the answer is B

OpenStudy (anonymous):

Thank you!!

OpenStudy (anonymous):

welcome